If f(x)= 1/3x^3 + 1/3x^2 + 1/3x + 1/3 and g(x)= x^2 + 1/3x + 1
calculate final answers as rational (fraction) numbers reduced to the lowest terms.
PLEASE HELP...THIS PROBLEM IS VERY CONFUSING. I WOULD APPRECIATE IT A LOT!

Answer by stanbon(75887)

If f(x)= 1/3x^3 + 1/3x^2 + 1/3x + 1/3 and g(x)= x^2 + 1/3x + 1
calculate final answers as rational (fraction) numbers reduced to the lowest terms.
------------------------
f(x)= 1/3x^3 + 1/3x^2 + 1/3x + 1/3
If I assume the x-factors are in the denominator I get:
A least common denominator of 2x^3.
Then f(x)=1/lcd +x/lcd + x^2/lcd +x^3/lcd
Rewrite to get:
f(x)=[1+x+x^2+x^3]/3x^3
-----------------
g(x)= x^2 + 1/3x + 1
lcd=3x
g(x)= 3x^3/lcd + 1/lcd + 3x/lcd
g(x)=[3x^3+1+3x]/3x
----------
Cheers,
Stan H.